# Sequential FDR and pFDR control under arbitrary dependence, with application to pharmacovigilance database monitoring

### v1

@@ 7 Conclusion
We have proposed both open-ended and truncated sequential multiple testing procedures which can control FDR or pFDR (and their type II analogs, FNR or pFNR in the open-ended case) under arbitrary dependence between the data streams. These procedures have shown large savings in average sample size compared to their fixed-sample equivalents in our simulation studies, and in Section 6 we demonstrate how these procedures may be used to monitor data streams like those coming from a pharmacovigilance database, like the UK’s Yellow Card database.
@@
Guo and Rao’s (2008, expression (19)) second constraint involves bounding from above the probability that a $p$-value for a true null falls in the interval $[\alpha_{j-1},\alpha_{j})$ by $\alpha_{j}-\alpha_{j-1}$. Validity of the $p$-value is not sufficient for this to hold and, for example, it may fail for valid but discrete $p$-values. Therefore their proof actually requires a condition stronger than validity on the $p$-values, for example having the exact uniform $(0,1)$ null distribution. We avoid the need for such a stronger condition by summing over $j$ in (47), and thus only need (10) which is the analog of validity for $p$-values in the fixed sample set up.

### v2

@@ 7 Conclusion
We have proposed both open-ended and truncated sequential multiple testing procedures which can control FDR or pFDR (and their type II analogs, FNR or pFNR in the open-ended case) under arbitrary dependence between the data streams. These procedures have shown large savings in average sample size compared to their fixed-sample counterparts in our simulation studies, and in Section 6 we demonstrate how these procedures may be used to monitor data streams like those coming from a pharmacovigilance database, like the UK’s Yellow Card database.
@@
Guo and Rao’s (2008, expression (19)) second constraint involves bounding from above the probability that a $p$-value for a true null falls in the interval $[\alpha_{j-1},\alpha_{j})$ by $\alpha_{j}-\alpha_{j-1}$. Validity of the $p$-value is not sufficient for this to hold and, for example, it may fail for valid but discrete $p$-values. Therefore their proof actually requires a condition stronger than validity on the $p$-values, for example having the exact uniform $(0,1)$ null distribution. We avoid the need for such a stronger condition by summing over $j$ in (47), and thus only need (10) which is the sequential analog of validity for $p$-values in the fixed sample set up.